A ski jumper speed down a slope and exits the ski track in a horizontal direction with a velocity of 25 m/s where the landing inclines falls off with a slope of 33 degrees.

1. What is the halftime of flight of the ski jump?

2. What is the range of the ski jumper when it lands?



1
Expert's answer
2021-09-15T11:38:03-0400

Given:

"v_0=25\:\rm m/s"

"\theta=33^{\circ}"


1. the halftime of flight

"t=\frac{v_0\sin\theta}{g}=\frac{25\sin33^{\circ}}{9.8}=1.4\:\rm s"

2. the range

"R=\frac{v_0^2\sin2\theta}{g}=\frac{25^2\sin66^{\circ}}{9.8}=58\:\rm m"